the diameter of a deoxyribonucleic acid (DNA) molecule is about 20 angstroms.express the surface area of the molecule in square meters and square inches.
1
Expert's answer
2013-04-26T10:01:22-0400
Idealizing we'll assume the molecule is sphere-shaped.
Surface area of the sphere is S = 4*pi*r^2 = 4*pi*(d/2)^2 =pi*d^2.
If d = 20 angstroms = 20 * 10^(-10) m = 2 * 10^(-9) m;
S = pi * (2 * 10^(-9) m)^2 = 1.2566 * 10^(-17) m^2
Let's also convert the diameter into inches d = 2 * 10^(-9)m * (1 inch / .0254 m) = 7.874 * 10^(-8) inch;
S = pi * ( 7.874 * 10^(-8) inch)^2 = 1.948 * 10^(-14)inch^2.